OpenStudy (anonymous):

What is the slope of a roof that rises 12 feet over a run of 28 feet?

OpenStudy (anonymous):

slope = rise / run this should be the information you need

OpenStudy (studygurl14):

\[\frac{12}{28}\rightarrow \frac{3}{7}\]

OpenStudy (studygurl14):

slope = 3/7